One leg of a right triangle is increased by 40% and the other leg is increased by 15%. By what percent is the area of the new righttriangle increased?
Answer by stanbon(48545)

One leg of a right triangle is increased by 40% and the other leg is increased by 15%. By what percent is the area of the new righttriangle increased?
--------------
Original Rectangle:
Area = (1/2)bh
---------
Enlarged Rectangle:
Area = (1/2)*(1.4b)(1.15h)
Area = (1.4*1.15)(1/2 bh)
Area = 1.61 (1/2 bh)
----------------
The Area has increased by 61%
Cheers,
Stan H.